# Bayesian brain mapping: population-informed individualized functional topography and connectivity

### v1

@@ 2.3 Fitting the BBM model to single-subject data
Once the priors have been estimated, the BBM model can be fit for a single individual, using a single or multiple fMRI sessions, to estimate topography and connectivity. Model fitting is implemented in the BayesBrainMap function BrainMap() through an iterative procedure based on expectation-maximization (Mejia et al.,, 2020, 2023) or variational Bayes (Mejia et al.,, 2025), depending on the specific model. This function produces point estimates (i.e., posterior means) of the subject-specific spatial maps and FC matrices, as well as measures of uncertainty (i.e., posterior variances) to facilitate inference.
@@
Finally, areas of significant engagement for each network are identified based on posterior inference using the function engagements(). Because the BBM framework has relatively high power to detect non-zero engagements, we can specify a minimum effect size of interest for inference. Since the spatial maps are in arbitrary units, we draw inspiration from a common method for thresholding group ICA maps, where values over $z$ standard deviations above the mean are commonly displayed, while values below this are considered negligible and suppressed. Following this logic, we set a minimum effect size for inference as the prior mean value corresponding to $z$ standard deviations above the mean. The engagements() function allows the user to specify $z$, or to provide multiple values for $z$ to produce a nested set of significant engagements of different strengths (illustrated in Figure 5).
@@
In the Github demo that accompanies this manuscript, we illustrate the use of the BrainMap() and engagements() functions to analyze several individual participants from the HCP.
@@ 3 Results
Spatial topography prior means for several components from the Yeo17 template are shown in Figure 2. While the template is constrained to have no overlap between networks, the prior relaxes that constraint. The patterns seen in the prior mean maps reflects the engagement of the network in the training sample, with no constraint to encourage or enforce spatial independence. This additional flexibility leads to expansion of the template networks beyond their original boundaries, suggesting the presence of regions that engage in multiple networks as well as population heterogeneity in the precise localization of networks.
Arrows indicate areas that differ markedly between the parcels and the prior mean maps. For instance, in the Default A network, bilateral, roughly symmetric temporal lobe engagement is seen in the prior, whereas the parcel only includes temporal lobe engagement within the right hemisphere. Another noteworthy example is engagement of Broca’s area and other areas commonly associated with language in the temporal parietal network, while the corresponding parcel is constrained to the temporal lobe. Similar patterns are often seen in networks characterized by strong engagement in the superior temporal gyrus and the temporal parietal junction when using ICA or other network estimation methods that allow for overlap between networks.
@@
The pairwise spatial overlap between networks for different templates and their corresponding priors is shown in Appendix Figure 6. Overlap is quantified using the Dice coefficient. For continuous network maps, we first threshold the maps at a z-score of $\pm 2$, following a common practice for thresholding group ICA maps to isolate areas of engagement. The parcellation-based templates by definition have zero overlap between networks, but their corresponding priors exhibit overlap between certain networks. Network maps based on PROFUMO exhibit the most overlap, reflecting a key feature of PROFUMO, and the degree of overlap is similar in the template and the prior. To a slightly lesser degree, the GICA template and prior also exhibit overlap, contradicting a common belief that ICA does not allow for spatial overlap. Spatial ICA components are optimized to maximize statistical independence, which does not imply non-overlap. On the contrary, non-overlap implies anti-correlation (or from probability, disjoint events are never independent).
@@
Figure 3: HCP-derived spatial topography priors for a default mode network component from various templates. Templates include two parcellations (Yeo17 and MSC) and two types of network maps (ICA and PROFUMO) and are ordered from lowest to highest resolution (number of networks). The specific networks displayed are PROFUMO mode 8, GICA15 component 2, Yeo17 Deafult A, MSC Default, and GICA25 IC 2. Template maps are displayed on an arbitrary scale, while the scale of the priors is inversely proportional to the template resolution $Q$, given the additive nature of the BBM decomposition.
Spatial topography priors for a default mode network component from various templates are shown in Figure 3. Similar features are seen across the different priors, even though the templates are based on distinct model assumptions. The prior standard deviation maps show that individual variation in engagement is greatest in areas of high average engagement, reflecting differences in the configuration, size, and distribution of networks across individuals. The low prior variance in background areas, where true population variability is low, will result in a high degree of regularization in these areas for the corresponding networks. Higher prior variance in areas of engagement, by contrast, will allow for individual features to be expressed. Furthermore, due to the additive nature of the decomposition, noise reduction in background areas of certain networks will indirectly improve estimation of engagement in other networks corresponding to the same physical location.
@@
Figure 4: Example of the functional connectivity (FC) prior in BBM, based on the Yeo17 template. The element-wise empirical mean and standard deviation (SD) within the training set used to establish the priors are show in the first column. The second and third columns show the element-wise mean and SD based on the Cholesky and the inverse Wishart choices of prior. The population mean is captured by both priors, but only the Cholesky prior captures the population variance patterns.
The functional connectivity priors in BBM are illustrated in Figure 4 for the Yeo17 template and in Appendix Figure 7 for other templates. The element-wise empirical mean and standard deviation within the training set, representing population tendencies and variability, are shown in the first column. The second two columns show the corresponding values based on the two choices of FC prior in BBM: the permuted Cholesky prior described in Mejia et al., (2025) and the conjugate inverse Wishart (IW). While both priors capture the population mean, only the Cholesky prior accurately captures the population variance. This is because the IW distribution for covariance matrices includes a single scalar parameter that moderates the variance across the matrix, so that the element-wise variance is monotonically related to the element-wise mean. In BBM, this parameter is chosen to minimize the difference between the population and prior element-wise variance, but with a constraint that the prior variance should not be less than the population variance anywhere. This avoids an overly informative prior, which would result in too much shrinkage towards the group and loss of relevant individual information. Due to this limitation of the inverse Wishart distribution, the IW FC prior tends to have much higher variance for many FC edges compared with the population, resulting in weak shrinkage. The Cholesky prior, by contrast, is more informative while not being overly informative, which serves to maximize accuracy of FC estimates, and has indeed been shown to produce more reliable and individualized estimates of between-network FC (Mejia et al.,, 2025). Improved estimation of the temporal component of the BBM model may also have downstream benefits for estimation of spatial topography, since the temporal and spatial components are estimated iteratively.
@@
Figure 5: Example individual BBM spatial engagement maps. A single HCP subject was analyzed using BBM with population-derived priors based on the Yeo17 template. The BBM posterior mean and standard deviation for a network corresponding to the Yeo17 DefaultA parcel are shown, along with areas of statistically significant engagement. Significance is based on the Bayesian equivalent of a hypothesis test with $\alpha=0.05$ and Bonferroni correction. A range of effect size thresholds are specified based on $\sigma$, the standard deviation of the prior mean map. This is analogous to the common practice of thresholding group ICA maps at a certain number of standard deviations from the mean to isolate the main areas of engagement.
Figure 5 shows example individual-level spatial topography maps produced by applying BBM to analyze data from one HCP subject, using population-derived priors based on the Yeo17 template. A single network is shown corresponding to the DefaultA parcel. The posterior mean provides an estimate of engagement in the network at every vertex, while the posterior standard deviation represents uncertainty around that estimate. Binary maps of statistically significant engagement are produced based on the posterior mean and standard deviation, using the Bayesian equivalent of a hypothesis test. A range of minimum effect sizes are tested against, resulting in nested sets of significantly engaged vertices. The ability to specify a minimum effect size gives the researcher flexibility to focus on locations that exhibit intense network engagement, or to consider all vertices that exhibit engagement that is significantly different from zero. The researcher can also consider a range of minimum effect sizes, as we have done here, to distinguish between levels of engagement.
@@ 4 Discussion
In this paper we illustrate Bayesian brain mapping (BBM), a source separation technique for fMRI that leverages population information through population-derived priors to reliably estimate individual-level network topography and functional connectivity. Those priors use established network maps, such as parcellations or group ICA maps, as templates, producing individual-specific analogues of known networks. The priors relax constraints of the template, allowing for overlap between networks and differential engagement within networks. The powerful noise reduction properties of the priors facilitate reliable estimation of individualized functional topography with moderate scan duration, mitigating the need for extensive scanning of individuals. We also provide resources to facilitate its implementation, including a demo illustrating the BayesBrainMap R package functions for prior estimation, model fitting, and inference, as well as HCP-derived priors based on a range of templates.
@@
Several areas of future work are planned to extend BBM to a wider variety of settings and introduce new capabilities. These include transfer learning to adapt priors to new populations without large training samples, algorithms for computationally efficient spatial modeling to enhance accuracy and power, inclusion of covariates and nuisance effects such as site and scanner, and mixture priors for heterogeneous populations. All extensions will be implemented in BayesBrainMap as they are developed.

### v2

@@ 2.3 Fitting the single-subject model
Once the priors have been estimated, the BBM model can be fit to one or more fMRI sessions from a single individual to estimate functional brain topography and FC. Model fitting is implemented in the BayesBrainMap function fit_BBM() through an iterative procedure based on expectation-maximization (Mejia et al.,, 2020, 2023) or variational Bayes (Mejia et al.,, 2025), depending on the specific model. This function produces point estimates (i.e., posterior means) of the subject-specific spatial maps and FC matrices, as well as measures of uncertainty (i.e., posterior variances or samples) to facilitate inference.
@@
Finally, areas of significant engagement for each network are identified using posterior inference with the function id_engagements(), which by default applies false discovery rate (FDR) multiple comparisons correction. Because the BBM framework has relatively high power to detect non-zero engagements, we can specify a minimum effect size of interest for inference. Since the spatial maps are in arbitrary units, we draw inspiration from a common method for thresholding group ICA maps, where values over $z$ standard deviations above the mean across space are commonly displayed, while values below this are considered negligible and suppressed. Following this logic, we set a minimum effect size for inference as the prior mean value corresponding to $z$ standard deviations above the mean. The id_engagements() function allows the user to specify $z$, or to provide multiple values for $z$ to produce a nested set of significant engagements of different strengths (illustrated in Figure 5).
@@
In the Github demo that accompanies this manuscript, we illustrate the use of the fit_BBM() and id_engagements() functions to analyze several individual participants from the HCP.
@@ 3 Results
3.1 Priors relax spatial constraints of templates and allow network overlap
Prior means on the spatial network topography for several networks from the Yeo17 template are shown in Figure 2. As a hard parcellation, the template is constrained to have no overlap between networks. By contrast, the prior reflects the engagement of the network in the training sample, with no constraint to encourage or enforce spatial independence. This additional flexibility allows the template networks to extend beyond their original boundaries, suggesting the presence of regions that engage in multiple networks and/or individual differences in network boundaries.
Arrows indicate areas that differ markedly between the template parcels and the prior mean maps. For instance, in the Default A network, bilateral, roughly symmetric temporal lobe engagement is seen in the prior, whereas the template parcel only includes right-hemisphere temporal lobe engagement. Another noteworthy example is engagement of Broca’s area and other areas commonly associated with language in the temporal parietal network, while the corresponding parcel is constrained to the temporal lobe. The prior mean for this network is similar to patterns often seen in networks characterized by strong engagement in the superior temporal gyrus and the temporal parietal junction when using ICA, PROFUMO, or other network estimation methods that allow for overlap between networks.
@@ 3.1 Priors relax spatial constraints of templates and allow network overlap
The pairwise spatial overlap between networks for different templates and their corresponding priors is shown in Appendix Figure S.1. Overlap is quantified using the Dice coefficient. For continuous network templates and prior means, we first threshold the maps at a z-score of $\pm 2$ (where the z-score is based on the mean and standard deviation of the prior mean map across all vertices), following a common practice for thresholding group ICA maps to isolate areas of engagement. The parcellation templates, by definition, have zero overlap between networks, yet their corresponding priors exhibit substantial overlap between certain networks. Network maps based on PROFUMO exhibit the most overlap, reflecting a key feature of that method, and the degree of overlap is similar in the template and the prior. To a slightly lesser degree, the GICA template and prior also exhibit overlap. This illustrates that spatial independence in ICA does not prohibit spatial overlap, as is sometimes assumed. [^2] On the contrary, non-overlap implies anti-correlation, so non-correlated spatial maps will tend do exhibit overlap. Consider this from a probability perspective: disjoint events are never independent, since the occurrence of one event precludes occurrence of the other. Still, the GICA prior mean maps exhibit somewhat stronger overlap than the GICA templates, suggesting a relaxation of the spatial constraints in the ICA model.
@@
Figure 3: HCP-derived spatial topography priors for a default mode network component from various templates. Templates include two parcellations (Yeo17 and MSC) and two types of network maps (ICA and PROFUMO) and are ordered from lowest to highest resolution (number of networks). The specific networks displayed are PROFUMO mode 8, GICA15 component 2, Yeo17 Default A, MSC Default, and GICA25 IC 2. Template maps are displayed on an arbitrary scale, while the scale of the priors is inversely proportional to the template resolution $Q$, given the additive nature of the BBM decomposition.
3.2 Different templates produce similar network topography priors
Spatial topography priors for a default mode network component from various templates are shown in Figure 3. Similar features are seen across the different priors, even though the templates are based on distinct model assumptions. The prior variance maps show that individual variation in engagement is greatest in areas of high average engagement, reflecting differences in the configuration, size, and distribution of networks across individuals. The low prior variance in background areas, where true population variability is low, will result in a high degree of regularization in these areas for the corresponding networks. Higher prior variance in areas of engagement, by contrast, will allow for individual features to be expressed. Furthermore, due to the additive nature of the decomposition, noise reduction in background areas of certain networks will indirectly improve estimation of engagement in other networks corresponding to the same physical location.
@@ 3.2 Different templates produce similar network topography priors
Figure 4: Example of the BBM functional connectivity (FC) prior. Results shown here correspond to the Yeo17 template; results for other templates are displayed in Appendix Figure S.2. The first column shows the element-wise empirical mean and standard deviation (SD) within the training set, which is used to establish the priors. The second and third columns show the element-wise mean and SD based on the Cholesky and the inverse-Wishart priors. The population mean is captured by both priors. However, the inverse-Wishart prior fails to capture the element-wise population variance patterns, given that distribution’s limited flexibility. The Cholesky prior, which we developed for BBM, accurately captures the element-wise population variance patterns.
3.2.1 BBM FC prior accurately captures population variance patterns
The FC priors in BBM are illustrated in Figure 4 for the Yeo17 template and in Appendix Figure S.2 for other templates. The element-wise empirical mean and standard deviation within the training set, representing population tendencies and variability, are shown in the first column. The second two columns show the corresponding values based on the two choices of FC prior in BBM: the permuted Cholesky prior described in Mejia et al., (2025) and the conjugate inverse-Wishart (IW), which is also used in PROFUMO (Harrison et al.,, 2020). While both priors capture the population mean, only the Cholesky prior accurately captures the population variance. This is because the IW distribution for covariance matrices includes a single scalar parameter that moderates the variance across the matrix, so that the element-wise variance is monotonically related to the element-wise mean. In BBM, this parameter is chosen to minimize the difference between the population and prior element-wise variance, but with a constraint that the prior variance should not be less than the population variance anywhere. This avoids an overly informative prior, which would result in too much shrinkage towards the group and loss of relevant individual information. Due to this limitation of the inverse-Wishart distribution, the IW FC prior tends to have much higher variance for many FC edges compared with the population, resulting in unnecessarily weak shrinkage. The Cholesky prior, by contrast, is more informative while not being overly informative. This theoretically maximizes accuracy of FC estimates, and has indeed been shown to produce more reliable and individualized estimates of between-network FC (Mejia et al.,, 2025). Improved estimation of the temporal component of the BBM model may also have downstream benefits for estimation of spatial topography, since the temporal and spatial components are estimated iteratively.
Prior Mean Prior SD
Refer to caption Refer to caption
Refer to caption Refer to caption
@@ 3.2.1 BBM FC prior accurately captures population variance patterns
Figure 5: Example individual BBM spatial engagement maps. A single HCP subject was analyzed using BBM with population-derived priors based on the Yeo17 template. The BBM posterior mean and standard deviation for a network corresponding to the Yeo17 DefaultA parcel are shown, along with areas of statistically significant engagement. Significance is based on the Bayesian equivalent of a hypothesis test with $\alpha=0.05$ and Bonferroni correction within each network. A range of effect size thresholds are specified based on $\sigma$, the standard deviation of the prior mean map. This is analogous to the common practice of thresholding group ICA maps at a certain number of standard deviations from the mean to isolate the main areas of engagement.
3.3 Illustration of BBM for personalized network topography and inference
Figure 5 shows example personalized functional network topography maps produced by applying BBM to analyze data from one HCP subject, using the population-derived priors based on the Yeo17 template. A single network is shown corresponding to the Default A parcel. The posterior mean provides an estimate of engagement in the network at every vertex, while the posterior standard deviation represents uncertainty around that estimate. Binary maps of statistically significant engagement are produced based on the posterior mean and standard deviation, using the Bayesian equivalent of a hypothesis test. A range of “null hypotheses” based on different minimum effect sizes of interest are tested against, resulting in nested sets of significantly engaged vertices. The ability to specify a minimum effect size gives the researcher flexibility to focus on locations that exhibit intense network engagement, or to consider all vertices that exhibit non-zero engagement. The researcher can also consider a range of minimum effect sizes, as we have done here, to distinguish between levels of engagement.
3.4 Computational demands of BBM
Prior estimation based on the templates listed in Table 1 was completed for the HCP population of 348 subjects in an average of 155 minutes per template, with peak memory usage of 126 GB. All calculations were performed using 48 concurrent threads on an Intel Xeon W 2.7 GHz processor running macOS Tahoe 26.2. The number of subjects in the sample, the number of volumes in each scan, and the number of vertices will directly affect the computational demands of estimate_prior().
It is worth noting that this calculation needs to be performed only once for each training dataset. For representative populations, priors can ideally be shared through a repository (OSF in the present study), thereby eliminating the need for repeated prior estimations and allowing users to proceed directly to the fit_BBM() function, which fits the model to individual data for the focal subject(s) of the study.
@@ 4 Discussion
In this paper we have described and illustrated Bayesian brain mapping (BBM), a personalized functional brain network mapping technique that leverages population information through population-derived priors for highly reliable network topography and between-network FC estimation. Those priors use established network maps or parcellations as templates, producing individual-specific analogues of known networks. The priors relax constraints of the template, allowing for overlap between networks and differential engagement within networks. The powerful noise reduction properties of the priors facilitate reliable estimation of individualized functional topography with moderate scan duration, mitigating the need for extensive scanning of individuals (Mejia et al.,, 2020; Derman,, 2026). We also provide resources to facilitate its implementation, including a demo illustrating the BayesBrainMap R package functions for prior estimation, model fitting, and inference, as well as HCP-derived priors based on a range of templates.
BBM lowers the barrier to personalized functional brain network mapping by reducing required scan duration through noise-mitigating informative priors, while avoiding heavy computational demands associated with multi-subject hierarchical models. Furthermore, once the priors have been established, model estimation is fast and parallelizable over multiple subjects. For experiments involving healthy adults, the HCP-derived priors we provide here can be applied directly, saving time and simplifying the analysis pipeline. For studies of other populations, our pipeline can be adapted to produce priors using other publicly available resources or data from the focal study.
Several limitations of the current BBM framework should be noted. First, priors should only be applied to study individuals coming from the same or similar populations. As such, the HCP-derived priors we share should not be utilized for the study of populations other than healthy young adults, such as children or the elderly. Other large publicly available datasets available facilitate the use of BBM for certain populations including Alzheimer’s disease (Mueller et al.,, 2005), adolescents (Casey et al.,, 2018), infants (Edwards et al.,, 2022), and many others. However, for studies of rarer populations it may be difficult to achieve samples large enough to construct accurate BBM priors with the current prior estimation methods. Second, BBM is currently designed for application to a single population; studies involving multiple groups typically construct priors based on equal representation from each group, but a mixture distribution would arguably better represent the population in this context. Third, BBM does not currently incorporate harmonization for multi-site datasets, which is a limitation for consortium datasets like the Alzheimer’s Disease Neuroimaging Initiative (Mueller et al.,, 2005) or the Autism Brain Imaging Data Exchange (Di Martino et al.,, 2014). Finally, while it is possible to incorporate spatial priors into BBM for added accuracy and power (Mejia et al.,, 2023), the current implementation is computationally intensive, limiting its practical utility.
@@
Several areas of future work are planned to address these limitations and extend the capabilities of BBM. These include transfer learning to adapt priors to new populations without large training samples, algorithms for computationally efficient spatial modeling to enhance accuracy and power, inclusion of covariates and nuisance effects such as site and scanner, and mixture priors for heterogeneous populations. All extensions will be implemented in the BayesBrainMap R package as they are developed. We also plan to derive and share topography and connectivity priors from other publicly available resources to facilitate the use of Bayesian brain mapping for personalized functional brain mapping in a wider variety of populations.
5 Conclusion
Bayesian brain mapping (BBM) provides a scalable, flexible, and computationally convenient framework for personalized functional brain network mapping. BBM employs population-derived priors corresponding to established network maps to mitigate noise, thus enhancing accuracy and power while reducing scan duration requirements. Several resources accompanying this manuscript, including the BayesBrainMap R package, pre-estimated HCP-derived priors, and user-friendly demos, lower the barrier to adoption of BBM for personalized functional network mapping. The BBM framework and these openly available resources serve to make precision functional mapping accessible across a wider range of research and clinical contexts.
6 Acknowledgments
This work was supported in part by the National Institutes of Health through the National Institute on Aging [Grant No. R01 AG083919 (to A.F.M.)], the National Institute of Mental Health’s National Research Service Award [Grant No. F31 MH140552 (to E.R.B.)], and the National Science Foundation Graduate Research Fellowship [Grant No. DGE-2234667 (to E.R.B.)].